# Efficient Speech Translation through Model Compression and Knowledge Distillation

### v1

@@ 4.4 Pruning more layers
In our main pruning experiments (cf. Section 3.3, we pruned only 8 layers based on layer importance evaluation. We decided to experiment with pruning more layers to explore the level to which a model can be pruned while keeping a similar level of translation performance. As Table 6 illustrates, we compare pruning 8, 10, 12, and 16 layers. We observe that the quality after pruning up to 12 layers and fine-tuning the pruned model on the ACL 60/60 dataset is close to pruning 8 layers. However, when pruning 16 layers, the quality starts to degrade considerably. It is worth noting that the results reported in Table 6 are only for the pruning and initial fine-tuning steps, while Section 3.3 describes the whole process that involves knowledge distillation, extra compression with quantization, and further fine-tuning with QLoRA.
@@ 4.5 Fine-tuning before/after pruning
It is common to fine-tune pruned models after pruning (Kim et al., 2023). As we pruned 1/4 of the decoder layers, obtaining valid translations for German was not possible without further training (cf. Table 2). Similarly, fine-tuning the teacher model on in-domain data before pruning is especially recommended for downstreaming tasks (Li et al., 2020).
@@ 4.6 Out-of-Domain Data Size
After layer pruning, we added more data for fine-tuning the pruned model to recover the translation quality of the teacher model. We mixed the in-domain data (ACL 60/60), knowledge distillation data, and out-of domain data from the CoVoST2 dataset. We experimented with different sizes of out-of-domain data, namely 10k, 50k, 80k and 100k randomly sampled segments. As Table 7 shows, for English-to-German translation, increasing the data size from 10k to 50k and 80k segments improves the overall performance. However, increasing the data size to 100k has diminishing returns, especially in terms of BLEU and chrF++ scores, while it improves the COMET score.
@@
Table 7: Investigating the effect of the data sizes used from the out-of-domain CoVoST2 dataset. Increasing out-of-domain data from 10k to 50k or 80k improves the translation quality on the German-to-English test split of the ACL 60/60 dataset. When increasing the out-of-domain data from 80k to 100k, this only improves the COMET score, but not the BLEU and chrF++ scores.
@@ 5 Inference and Evaluation
For inference, we use greedy generation by disabling the sampling options. We apply the prompt illustrated in Section 4.1. We use a batch size of 1, and set the generation max length of 1024 tokens.
To evaluate our systems, we calculated BLEU (Papineni et al., 2002), chrF++ (Popović, 2017), as implemented in the sacreBLEU library [^7] https://github.com/mjpost/sacrebleu (Post, 2018). While we use chrF++ for German, we only use the raw chrF for Chinese, following the author of the metric who noted that “the concept of Chinese words is generally not clear” (Popović, 2017). For semantic evaluation, we used COMET (Rei et al., 2020). [^8] In particular, we used the “wmt20-comet-da” model. Table 1 and Table 2 report the results of the main experiments. Moreover, we conducted an ablation study (cf. Section 4) to investigate the effect of modifying some aspects of our experiments, such as the baseline model, the pruning approach, the number of pruned layers, and the data size.
@@ 6 Conclusions and Future Work
In this work, we showed that combining multiple compression techniques enables substantial model size reduction with minimal impact on speech translation performance. To conclude, QLoRA fine-tuning with knowledge distillation achieved superior translation quality compared to layer pruning, though with reduced model compression. To achieve higher compression ratios while preserving translation quality, we employed a combined approach using iterative layer pruning, quantization, knowledge distillation, and multi-stage fine-tuning. The code of our experiments is publicly available. [^9] https://github.com/ymoslem/Model-Compression

### v2

@@ 4.5 Pruning more layers
In our main pruning experiments (cf. Section 3.3), we pruned only 8 layers based on layer importance evaluation. We decided to experiment with pruning more layers to explore the level to which a model can be pruned while keeping a similar level of translation performance. As Table 6 illustrates, we compare pruning 8, 10, 12, and 16 layers. We observe that the quality after pruning up to 12 layers and fine-tuning the pruned model on the ACL 60/60 dataset is close to pruning 8 layers. However, when pruning 16 layers, the quality starts to degrade considerably. It is worth noting that the results reported in Table 6 are only for the pruning and initial fine-tuning steps, while Section 3.3 describes the whole process that involves knowledge distillation, extra compression with quantization, and further fine-tuning with QLoRA.
Pruning reduces storage footprint while accelerating inference speed by approximately 20% and 40% when 8 and 16 layers are pruned, respectively, compared to the unpruned baseline, after full-parameter fine-tuning of both models on the English-to-German in-domain data. In contrast, 4-bit quantization as implemented in QLoRA reduces storage at the cost of inference speed. Given that the shared task prioritized minimizing storage requirements, we applied QLoRA to the pruned model in the final fine-tuning stage. For deployment scenarios where inference speed is critical, however, the pruned model can be fine-tuned with standard LoRA instead of QLoRA to avoid quantization overhead.
@@ 4.6 Fine-tuning before/after pruning
It is common to fine-tune pruned models after pruning (Kim et al., 2023). As we pruned 1/4 of the decoder layers, obtaining valid translations for German was not possible without further training (cf. Table 2). Similarly, fine-tuning the teacher model on in-domain data before pruning is especially recommended for downstream tasks (Li et al., 2020).
@@ 4.7 Out-of-Domain Data Size
After layer pruning, we added more data for fine-tuning the pruned model to recover the translation quality of the teacher model. We mixed the in-domain data (ACL 60/60), knowledge distillation data, and out-of-domain data from the CoVoST2 dataset. We experimented with different sizes of out-of-domain data, namely 10k, 50k, 80k and 100k randomly sampled segments.
@@
Table 8: Investigating the effect of the data sizes used from the out-of-domain CoVoST2 dataset. Increasing out-of-domain data from 10k to 50k or 80k improves the translation quality on the English-to-German test split of the ACL 60/60 dataset. When increasing the out-of-domain data from 80k to 100k, this only improves the COMET score, but not the BLEU and chrF++ scores.
As Table 8 shows, for English-to-German translation, increasing the data size from 10k to 50k and 80k segments improves the overall performance. However, increasing the data size to 100k has diminishing returns, especially in terms of BLEU and chrF++ scores, while it improves the COMET score.
@@ 5 Inference and Evaluation
For inference, we use greedy generation by disabling the sampling options. We apply the prompt illustrated in Section 4.1. We use a batch size of 1, and set the generation max length to 1024 tokens.
To evaluate our systems, we calculated BLEU (Papineni et al., 2002), chrF++ (Popović, 2017), as implemented in the sacreBLEU library [^7] https://github.com/mjpost/sacrebleu (Post, 2018). While we use chrF++ for German, we use raw chrF for Chinese, following the author of the metric who noted that “the concept of Chinese words is generally not clear” (Popović, 2017). For semantic evaluation, we use COMET (Rei et al., 2020). [^8] In particular, we used the “wmt20-comet-da” model. Table 1 and Table 2 report the results of the main experiments. Moreover, we conduct an ablation study (cf. Section 4) to investigate the effect of modifying some aspects of our experiments, such as the baseline model, the pruning approach, the number of pruned layers, and the data size.
@@ 6 Conclusions and Future Work
In this work, we showed that combining multiple compression techniques enables substantial model size reduction with minimal impact on speech translation performance. To conclude, QLoRA fine-tuning with knowledge distillation achieved superior translation quality compared to layer pruning alone, though with reduced model compression. To achieve higher compression ratios while preserving translation quality, we employed a combined approach using iterative layer pruning, quantization, knowledge distillation, and multi-stage fine-tuning. The code of our experiments is publicly available. [^9] https://github.com/ymoslem/Model-Compression